# Xiaobo Ma12, Pengyu Pan12, Jianfeng Li12, Wei Wang3, Weizhi Meng4, Xiaohong Guan12

### 10 Mar 2024 v1

@@ III-C Indistinguishable Transaction Feature
$\bullet$ Gaussian Distribution Fitting. Next, we employ the Gaussian distribution to model the fee distribution observed in real data. To achieve this, We use the set $K$ to denote all real data. We group transactions based on the inputCnt and outputCnt parameters, denoted as $K_{ij}$ for transactions with $i$ inputs and $j$ outputs.
@@
We further divide each transaction set into 10 intervals based on the percentile of inputsAmount. The $n$th interval in $K_{ij}$ can be represented as
$K_{ij}^{n}=\{T_{m}\in K_{ij}:A_{start_{n}}<inputsAmount_{m}\leq A_{end_{n}}\},$ (3)
where
$\footnotesize\begin{split}start_{n}&=p_{n}\times\lvert A\rvert,\\ end_{n}&=(p_{n}+0.1)\times\lvert A\rvert,\\ \forall p_{n}&\in P,P=\{0,0.1,0.2,\ldots,0.9\},\end{split}$ (4)
[…]
$inputsAmount_{m}\leq inputsAmount_{m+1})\}.$
We fit a Gaussian distribution to the fee distribution within each interval and represent it as $N_{ij}^{n}$.
$\footnotesize\begin{split}F_{ij}^{n}&=\{fee_{m}\in T_{m}:T_{m}\in K_{ij}^{n}\},\\ \mu&=\frac{1}{\lvert F_{ij}^{n}\rvert}\sum_{fee_{m}\in F_{ij}^{n}}fee_{m},\\ \sigma^{2}&=\frac{1}{\lvert F_{ij}^{n}\rvert}\sum_{fee_{m}\in F_{ij}^{n}}(fee_{m}-\mu)^{2},\\ N_{ij}^{n}&=N(\mu,\sigma^{2}).\\ \end{split}$ (6)
Additionally, we record the maximum and minimum fee values in $K_{ij}^{n}$ as $Max_{ij}^{n}$ and $Min_{ij}^{n}$, respectively. We use $Len_{ij}^{n}$ to denote the number of transactions in $K_{ij}^{n}$. To generate fake fee values, we sample $Len_{ij}^{n}$ data from $N_{ij}^{n}$ and remove data out of the range between $Max_{ij}^{n}$ and $Min_{ij}^{n}$. We denote the resulting fee set as $X_{ij}^{n}$.
$\footnotesize\forall x\in X_{ij}^{n},x\sim N_{ij}^{n}.$ (7)
Since the fee is a positive integer, appropriate rounding methods are applied to process the data in $X_{ij}^{n}$.
$\bullet$ Transaction Feature Sampling. Finally, we associate the three parameters of each covert transaction in the fake data with the corresponding $X_{ij}^{n}$. Random selection is employed to assign a fee value to each transaction. It is worth noting that in a transaction, the outputsAmount is obtained by subtracting the fee from the inputsAmount. With this step, the generation of the five parameters in the fake data is complete. To optimize the accuracy and realism of our fake data, our methodology extends beyond the mere use of GAN networks or Gaussian distributions. Instead of directly sampling real transaction parameters from the dataset, we adopt a more nuanced approach. Next, we elaborate on the rationale behind this approach.
Taking datasets BTD129 and BTD30 as examples (see Sec. IV-B for details), it is observed that inputsAmount and fee exhibit a wide range of values, with 99th percentiles of 15,666,023,540 and 31,671, respectively. Considering the sparse distribution resulting from the extensive range of inputsAmount in the datasets containing approximately 6.8 million transactions, relying solely on the GAN networks to generate all parameters does not yield satisfactory results. Additionally, using Gaussian distribution to fit each parameter independently overlooks the interdependencies between the parameters, as they are randomly combined. Directly sampling parameters from real data would result in many covert transactions having the same parameters as real transactions, drawing the attackers’ attention.

### 24 Mar 2024 v2

@@ III-C Indistinguishable Transaction Feature
$\bullet$ Discrete Distribution Fitting. Next, we perform a fitting analysis on the distribution of transaction fees within real data. To achieve this, We use the set $K$ to denote all real data. We group transactions based on the inputCnt and outputCnt parameters, denoted as $K_{ij}$ for transactions with $i$ inputs and $j$ outputs.
@@
We further divide each transaction set into n intervals based on the percentile of inputsAmount. We denote the collection of inputsAmount in the $n$th interval as $A^{n}$, with $A^{n}_{min}$ and $A^{n}_{max}$ representing the minimum and maximum values within this collection, respectively. The $n$th interval in $K_{ij}$ can be represented as
$K_{ij}^{n}=\{T_{m}\in K_{ij}:A^{n}_{min}<inputsAmount_{m}\leq A^{n}_{max}\}.$ (3)
The transaction fees within $K^{n}_{ij}$ can be represented as
$\footnotesize\begin{split}F_{ij}^{n}&=\{fee_{m}\in T_{m}:T_{m}\in K_{ij}^{n}\}.\end{split}$ (4)
We perform a discrete distribution fitting on $F^{n}_{ij}$ and generate $Len^{n}_{ij}$ transaction fees from it, denoted as $X^{n}_{ij}$. Here, $Len^{n}_{ij}$ is used to represent the number of transactions within $K^{n}_{ij}$.
$\bullet$ Transaction Feature Sampling. Finally, we associate the three parameters of each covert transaction in the fake data with the corresponding $X_{ij}^{n}$. Random selection is employed to assign a fee value to each transaction. It is worth noting that in a transaction, the outputsAmount is obtained by subtracting the fee from the inputsAmount. With this step, the generation of the five parameters in the fake data is complete. To optimize the accuracy and realism of our fake data, our methodology extends beyond the mere use of GAN or distribution fitting. Instead of directly sampling real transaction parameters from the dataset, we adopt a more nuanced approach. Next, we elaborate on the rationale behind this approach.
Taking datasets BTD129 and BTD30 as examples (see Sec. IV-B for details), it is observed that inputsAmount and fee exhibit a wide range of values, with 99th percentiles of 15,666,023,540 and 31,671, respectively. Considering the sparse distribution resulting from the extensive range of inputsAmount in the datasets containing approximately 6.8 million transactions, relying solely on the GAN to generate all parameters does not yield satisfactory results. The reason is that GAN struggles to effectively learn the relationships between various parameters. Similarly, we have not performed distribution fitting for each parameter individually. Directly sampling parameters from real data would result in many covert transactions having the same parameters as real transactions, drawing the attackers’ attention. In contrast to inputsAmount, the distribution of fee is not considered sparse. This implies that many transactions will have identical transaction fees. Therefore, we only fit and sample from the distribution of fee, aiming to achieve more realistic simulation of transaction parameters, while simultaneously reducing the attention of attackers.